OpenStudy (anonymous):

If one measurement of a golden rectangle is 8.8 inches, which could be the other measurement?

OpenStudy (mathstudent55):

The golden ratio is: \( 1 : \dfrac{1 + \sqrt{5}}{2} \)

OpenStudy (anonymous):

A golden rectangle has sides in proportion to the golden ratio. The golden ratio is known to be: \[\frac{ 1+\sqrt{5} }{ 2 }\approx1.618033\] In our rectangle we know that one measure is 8.8". This leads to two possible cases: 1. 8.8 is the long measure. 2. 8.8 is the short measure. Case 1: 8.8 is the short measure. The other measure is then: \[8.8*\frac{1+\sqrt{5}}{2} = \frac{8+8\sqrt{5}}{2}\approx14.2386\] Case 2: 8.8 is the short measure. The other side is then: \[8.8*\frac{2}{1+\sqrt{5}} = \frac{17.6}{1+\sqrt{5}}\approx5.4387\] The two possibilities for the other measurement are: 4.24 and 5.44

OpenStudy (mathstudent55):

^14.24

OpenStudy (anonymous):

Ah yes, my mistake. 14.32. Just a typo.

OpenStudy (mathstudent55):

\(1 : \dfrac{1 + \sqrt{5}}{2} \) is approximately the same as \(1 : 1.61804} 8.8 * 1.61804 = 14.24 8.8/1.61804 = 5.43